A 0.29-kg stone is held 1.5 m above the top edge of a water well and then dropped into it. The well has a depth of 5.5 m.

(a) Relative to the configuration with the stone at the top edge of the well, what is the gravitational potential energy of the stone−Earth system before the stone is released?
J

(b) Relative to the configuration with the stone at the top edge of the well, what is the gravitational potential energy of the stone−Earth system when it reaches the bottom of the well?
J

(c) What is the change in gravitational potential energy of the system from release to reaching the bottom of the well?
J
Answer #1

given
m = 0.29 kg
h = 1.5 m
d = 5.5 m

a) Relative to the configuration with the stone at the top edge of the well, the gravitational potential energy of the stone−Earth system before the stone is released

Ui = m*g*h

= 0.29*9.8*1.5

= 4.263 J <<<<<<<<------------Answer

b) Relative to the configuration with the stone at the top edge of the well, the gravitational potential energy of the stone−Earth system when it reaches the bottom of the well

Uf = m*g*(-d)

= 0.29*9.8*(-5.5)

= -15.631 J <<<<<<<<------------Answer

c) the change in gravitational potential energy of the system from release to reaching the bottom of the well,

Uf - Ui = -15.631 - 4.263

= -19.894 J <<<<<<<<------------Answer